OpenStudy (anonymous):

Find the area of the region bounded by the graphs of y = x^2; x = 0; x = 2; and the x axis.

OpenStudy (anonymous):

integrate ion of x^2 =(x^3)/3 thus the answer is 8/3 -0/3 =8/3

OpenStudy (anonymous):

the area you need is \[\begin{align} \int_{0}^{2}x^{2}dx &=\frac{x^{3}}{3} \text{ from 0 to 2. }\\ &=\frac{2^{3}}{3}-\frac{0^{3}}{3}\\ &=\: \frac{8}{3}\\ \end{align}\]